# Blameless Users in a Clean Room: Defining Copyright Protection for Generative Models

### v1

@@ Abstract
Are there any conditions under which a generative model’s outputs are guaranteed not to infringe the copyrights of its training data? This is the question of “provable copyright protection” first posed by Vyas, Kakade, and Barak [VKB23]. They define near access-freeness (NAF) and propose it as sufficient for protection. This paper revisits the question and establishes new foundations for provable copyright protection—foundations that are firmer both technically and legally. First, we show that NAF alone does not prevent infringement. In fact, NAF models can enable verbatim copying, a blatant failure of copy protection that we dub being tainted. Then, we introduce our blameless copy protection framework for defining meaningful guarantees, and instantiate it with clean-room copy protection. Clean-room copy protection allows a user to control their risk of copying by behaving in a way that is unlikely to copy in a counterfactual “clean-room setting.” Finally, we formalize a common intuition about differential privacy and copyright by proving that DP implies clean-room copy protection when the dataset is golden, a copyright deduplication requirement.
@@ 1 Introduction
Ideally, the generative model would never reproduce copyrighted work. That’s unrealistic. Typical models can be prompted to generate copyrighted output: print the following foob [VKB23, LCG24]. It’s also unnecessary. In the example, copying is clearly the user’s fault.
@@
Our goal is to formalize such a guarantee—in a mathematically and legally rigorous way—and to study its feasibility. We’re after “provable copyright protection” at deployment time, first studied by Vyas, Kakade, and Barak [VKB23]. They propose near access-freeness (NAF) as an answer.
@@
    We prove that near access-free (NAF) models can enable verbatim copying. NAF is the first (and only other) attempt at a mathematical definition intended to offer “provable copyright protection” for generative models [VKB23]. The proof leverages NAF’s lack of protection against multiple prompts or data-dependent prompts.
@@
    We define clean-room copy protection (($\kappa,\beta)$-clean), a first instantiation of our framework (Section 7), drawing inspiration from clean-room design. A training algorithm is $(\kappa,\beta)$-clean if, for every user who copies in a (counterfactual) “clean-room environment” with probability $\leq\beta$, the probability of copying in the real world is $\leq\kappa$. Clean-room copy protection excludes tainted models under mild assumptions.
@@
    We prove that differential privacy (DP) implies clean-room copy protection, formalizing a common intuition about DP and copyright (Section 8). This holds when the dataset is golden, a copyright deduplication requirement. Thus, DP provides a way to bring copyrighted expression “into the clean room” without tainting the model.
@@
This paper can be read linearly from beginning to end. Depending on your interests, you might consider reading out of order at first, referencing earlier sections as needed. Sections 1.12, and 4 are required for anything that follows them. For near access-freeness and its limitations: read Section 3, Appendix B, and Section 5. For our approach to defining meaningful copyright protection: read Sections 57. For the protection offered by differential privacy: start with Section 8 for the high level claims, then read Section 7 to understand the precise formalization. For discussion of practical considerations: read Section 9.
@@
There is a lot of recent work on copyright questions for generative AI [CLG+23, LCG24, HLJ+23], and the first generation of cases are working their way through the courts (e.g., New York Times v Microsoft). But most are only tangentially related to goal of stating and proving formal guarantees against copyright infringement.
Vyas, Kakade, and Barak were the first to propose a mathematical property—near access-freeness (NAF)—aimed at “preventing deployment-time copyright infringement” [VKB23]. Elkin-Koren, Hacohen, Livni, and Moran argue that copyright cannot be “reduced to privacy” [EKHLM24]. Referring to both NAF and DP by umbrella term “algorithmic stability”, they argue that the sort of provable guarantees that [VKB23] and this paper seek do not capture copyright’s complexities. Li, Shen, and Kawaguchi propose and empirically evaluate an attack called VA3 against NAF [LSK24]. They provide good evidence that the $\mathsf{CP\text{-}k}$ algorithm of [VKB23] may not prevent infringement, but some uncertainty remains (see Appendix B.3). We discuss these three papers at length. There is also work on new NAF algorithms [GAZ+24, CKOX24].
Scheffler, Tromer, and Varia [STV22] give a complexity-theoretic account of substantial similarity, and a procedure for adjudicating disputes. In contrast, we treat substantial similarity as a black box. That differential privacy might protect against infringement has been suggested in [BLM20, HLJ+23, VKB23, EKHLM24, CKOX24].
@@ 2 The legal-technical interface
To use legal concepts within a mathematical formalism, we assume some exogenously-defined functions reflecting and operationalizing those concepts. They are the interface between math and law, enabling mathematical study of legal concepts without formalizing the concepts themselves.  Precise definitions of these concepts are out of scope (perhaps impossible, though see [STV22] for an attempt at formalizing substantial similarity).Three legal concepts are needed for this paper: substantial similarity, ideas, and access. This section defines functions $\mathsf{SubSim}$ and $\mathsf{ideas}$ for the first two. Section 7.1 operationalizes access.
The owner of a copyright in a work has exclusive rights to reproduce, distribute, display, adapt, and perform the work. Original works of authorship are eligible for copyright protections as soon as they are fixed in a tangible medium, and eventually expires. While the work is protected, it is in-copyright. Many types of works are eligible, including books, music, poetry, plays, choreography, photographs, video, and paintings. Copying without permission is sometimes permitted, including under the fair use exception [HLJ+23].
@@
Copyright does not protect ideas, only original expression. For example, the text and images in a cookbook may be copyrighted, but not the method of cooking a dish. To be afforded protection, there must be a minimum of creativity beyond the ideas. The function $\mathsf{ideas}$ defines the non-copyrightable ideas of a work, as opposed their expression.
@@
We never actually compute $\mathsf{SubSim}$ or $\mathsf{ideas}$. One might view them as capturing the “true” legal concept. Or one may take a more practical view. For example, interpreting $\mathsf{SubSim}(w)$ as reflecting the opinion of a typical judge or jury. However the reader wishes to interpret these functions is fine (up to any assumptions stated where used). In particular, our results hold even for $\mathsf{SubSim}(w)=\{w\}$ and $\mathsf{ideas}(w)=\bot$. In contrast, a model provider (or data provider) would need to actually work with the copyright dependency graph in Section 7. We discuss this limitation in Section 9.
@@ 3 Near access-free models do not provide provable copyright protection
Near access-freeness (NAF) is a mathematical definition for “provable copyright protection” [VKB23]. This section describes near access-freeness and its limitations. We prove that models can enable verbatim copying while still satisfying NAF. While NAF provides protection against a single prompt that is independent of the training data, it makes no guarantees against many prompts (composition) [LSK24], nor a single prompt derived from non-copyrightable ideas (not expression). We emphasize that our focus is definitions for provable copyright protection, where NAF falls short. Practically, NAF may still provide useful protection in many natural settings. For space, we defer algorithms, proofs, and additional discussion to Appendix B.
@@ 3.1 NAF Background
Near access-freeness (NAF) is the first (and only other) attempt at a mathematical definition intended to offer “provable copyright protection” for generative models [VKB23]. Motivated by the legal requirement of access, NAF is meant to show “that the defendant’s work is close to a work which was produced without access to the plaintiff’s work” [VKB23]. NAF is closely related to DP. But surprisingly, any training algorithm $\mathsf{Train}$ can be used in a black-box way to construct a NAF model without the extra noise that is the hallmark of DP (Theorem 3.3).
@@
Prior works discuss and critique the NAF framework in three ways: questioning NAF’s technical effectiveness [LSK24], its legal applicability [LCG24, EKHLM24], and its real-world practicality [HLJ+23, EKHLM24, LCG24, CKOX24].
$\ast$ Technically, Li, Shen, and Kawaguchi propose and empirically evaluate an attack called VA3 against NAF [LSK24]. See Appendix B.3 for a full discussion. They show that a black-box attacker, can reliably induce a model to produce outputs that infringe on a target work $c^{*}$. They also give a white-box prompt writing algorithm for diffusion models that greatly improves performance. VA3 provides good evidence that the $\mathsf{CP\text{-}k}$ algorithm of [VKB23] may not prevent infringement. Some uncertainty remains because the algorithm implemented in VA3 deviates from the original. The paper leaves open whether $k$-NAF (with fixed $k$) prevents copyright infringement, formalizes a flawed attack model, and avoids the underlying definitional questions almost entirely.
$\ast$ Legally, Elkin-Koren et al. [EKHLM24] give the most extensive response to NAF. They correctly argue that copyright cannot be “reduced to privacy.” However, this is a straw man of version of [VKB23] and the present paper (see Appendix A). The sharpest criticism is by Lee, Cooper, and Grimmelman who argue that NAF is simply wrong on the law [LCG24]. “It is not a defense to copyright infringement that you would have copied the work from somewhere else if you hadn’t copied it from the plaintiff.” See Appendix B.4 for more discussion of NAF’s legal relevance.
$\ast$ Practically, the main concern is that deduplicating data in the way needed to make NAF effective is infeasible [EKHLM24, LCG24, HLJ+23]. It is unrealistic to produce enough clean training data to pretrain foundation models, advances in deduplication notwithstanding. Moreover, NAF learning algorithms are computationally expensive and may lag in performance [CKOX24]. These concerns are justified and also apply to DP. Still, there may be settings where golden data is practical (Sec. 9), to say nothing of the value of theory.
@@ 3.2 NAF definition and the Copy Protection ($\mathsf{CP}$) algorithm
Near access-freeness (NAF) is defined with respect to a function $\mathsf{safe}$. The $\mathsf{safe}$ function maps a copyrighted data point $c\in\mathcal{C}$ to a generative model $\mathsf{safe}_{c}\in\mathcal{P}$ trained without access to $c$. An example $\mathsf{sharded\mbox{-}safe}$ is in Appendix B (Alg. 2). Thm. 3.3 presents the main feasibility result: any training algorithm $\mathsf{Train}$ can be used as a black-box to construct an NAF model $\mathsf{CP}$, short for copy protection.
@@
Theorem 3.3 (Copy Protection algorithm [VKB23]).
@@ 3.3.1 $\mathsf{CP}$ can regurgitate training data
We show that $\mathsf{CP}$—the main NAF algorithm of [VKB23]—can fail to protect against copying. Using a prompt containing no copyrightable expression, a user can cause the NAF model returned by $\mathsf{CP}$ to regurgitate copyrighted training data. [^3] This doesn’t violate Theorem 3.3. We use the fact that the $\mathsf{CP}$ algorithm is not actually $k$-NAF for any fixed $k$. It is $k_{x}$-NAF, where $k_{x}$ depends on the prompt $x$. Applied to our construction, the bound on $k_{x}$ given by Theorem 3.3 is not meaningful for $x\in\mathsf{ideas}(D)$. Adapting our construction to the alternate NAF algorithm $\mathsf{CP\text{-}k}$ in [VKB23] results in a model sometimes fails to terminate. This is based on an observation of Thomas Steinke [Ste].
Observe that $\mathsf{CP}$ is a black-box transformation from an underlying training algorithm $\mathsf{Train}$. Thus, $\mathsf{CP}$ is really a family of algorithms—one per $\mathsf{Train}$.
The following theorem states that there exists $\mathsf{Train}^{*}$ such that the resulting NAF algorithm $\mathsf{CP}^{*}$ fails in the following way. On prompt $x=\mathsf{ideas}(c)$ reproduces training datum $c$ verbatim.
@@
We defer the proof of (a slight generalization) of this theorem to Appendix B.2. The proof uses the tainted training algorithm $\mathsf{Train}^{*}=\mathsf{Train}_{\mathsf{kv}}$ (Algorithm 1) described in Example 5.2.

### v2

@@ Abstract
Are there any conditions under which a generative model’s outputs are guaranteed not to infringe the copyrights of its training data? This is the question of “provable copyright protection” first posed in [VyasKB23]. They define near access-freeness (NAF) and propose it as sufficient for protection. This paper revisits the question and establishes new foundations for provable copyright protection—foundations that are firmer both technically and legally. First, we show that NAF alone does not prevent infringement. In fact, NAF models can enable verbatim copying, a blatant failure of copy protection that we dub being tainted. Then, we introduce our blameless copy protection framework for defining meaningful guarantees, and instantiate it with clean-room copy protection. Clean-room copy protection allows a user to control their risk of copying by behaving in a way that is unlikely to copy in a counterfactual “clean-room setting.” Finally, we formalize a common intuition about differential privacy and copyright by proving that DP implies clean-room copy protection when the dataset is golden, a copyright deduplication requirement.
@@ 1 Introduction
Ideally, the generative model would never reproduce copyrighted work. That’s unrealistic. Typical models can be prompted to generate copyrighted output: print the following foob [VyasKB23, lee2024talkin]. It’s also unnecessary. In the example, the copying is clearly the user’s fault.
@@
Our goal is to formalize such a guarantee—in a mathematically and legally rigorous way—and to study its feasibility. We’re after “provable copyright protection” at deployment time, first studied by Vyas, Kakade, and Barak [VyasKB23]. They propose near access-freeness (NAF) as an answer.
@@
    We prove that near access-free (NAF) models can enable verbatim copying. NAF is the first (and only other) attempt at a mathematical definition intended to offer “provable copyright protection” for generative models [VyasKB23]. The proof leverages NAF’s lack of protection against multiple prompts or data-dependent prompts.
@@
    We define clean-room copy protection $($$\kappa,\beta)$-clean), a first instantiation of our framework (Section 7), drawing inspiration from clean-room design. A training algorithm is $(\kappa,\beta)$-clean if, for every user who copies in a (counterfactual) “clean-room environment” with probability $\leq\beta$, the probability of copying in the real world is $\leq\kappa$. Clean-room copy protection lets users choose their tolerance for risk $\kappa$ and then tune $\beta$ accordingly. Clean-room copy protection also excludes tainted models under mild assumptions.
@@
    We prove that differentially private (DP) training implies clean-room copy protection, formalizing a common intuition about DP and copyright (Section 8). This holds when the dataset is golden, a copyright deduplication requirement. Thus, DP provides a way to bring copyrighted expression “into the clean room” without tainting the model.
This paper does not offer a legal analysis of where liability would lie under existing law. Copyright infringement is generally treated as a strict liability tort, suggesting that the user may always liable. This paper is premised on the possibility that strict liability may sometimes be unjust when generative AI is involved. The simple fact is that there is no feasible, general way for a user to determine with confidence whether a particular model output infringes or not. Some legal scholars agree [chatterjee2019minds, goodyear2025artificial], reasoning from caselaw requiring “some element of volition or causation.” [^1] Religious Tech. Center v. Netcom On-Line Comm., 907 F. Supp. 1361 (N.D. Cal. 1995)
Instead of advancing such legal arguments, our paper develops a mathematical framework for preventing this sort of infringement—where holding the user liable would be unjust—and proposes an indemnification policy for assigning fault when copying does inevitably occur (Section 9).
@@
This paper can be read linearly from beginning to end. Depending on your interests, you might consider reading out of order at first, referencing earlier sections as needed. Sections 1.12, and 4 are required for anything that follows them. For near access-freeness and its limitations: read Sections 35. For our approach to defining meaningful copyright protection: read Sections 57. For the protection offered by differential privacy: start with Section 8 for the high level claims, then read Section 7 to understand the precise formalization. For discussion of practical and policy considerations: read Section 9.
@@
There is a lot of recent work on copyright questions for generative AI [cooper2023report, lee2024talkin, henderson2023foundation], and the first generation of cases are working their way through the courts (e.g., New York Times v Microsoft). But most are only tangentially related to goal of stating and proving formal guarantees against copyright infringement.
Vyas, Kakade, and Barak were the first to propose a mathematical property—near access-freeness (NAF)—aimed at “preventing deployment-time copyright infringement” [VyasKB23]. Elkin-Koren, Hacohen, Livni, and Moran argue that copyright cannot be “reduced to privacy” [elkinkoren2024copyright]. Referring to both NAF and DP by umbrella term “algorithmic stability”, they argue that the sort of provable guarantees that [VyasKB23] and this paper seek do not capture copyright’s complexities. Li, Shen, and Kawaguchi propose and empirically evaluate an attack called VA3 against NAF [li2024va3]. They provide good evidence that the $\mathsf{CP\text{-}k}$ algorithm of [VyasKB23] may not prevent infringement, but some uncertainty remains (see Appendix B.3). We discuss these three papers at length. There is also work on new NAF algorithms [golatkar2024cpr, chen2024randomization].
Scheffler, Tromer, and Varia [scheffler2022formalizing] give a complexity-theoretic account of substantial similarity, and a procedure for adjudicating disputes. In contrast, we treat substantial similarity as a black box. That differential privacy might protect against infringement has been suggested in [bousquet2020synthetic, henderson2023foundation, VyasKB23, elkinkoren2024copyright, chen2024randomization].
Livni, Moran, Nissim, and Pabbaraju [livni2024credit] study the problem of attributing credit to inputs of an algorithm that influence its output, motivated by copyright. They define a condition under which a given input need not be credited (with credit required otherwise). Under their definition, a differentially private algorithm never has to credit its inputs—a manifestation of the common intuition that differential privacy prevents copying which we discuss in Section 8.
@@ 1.1 Notation
Symbol Definition
$\mathcal{W};~w,y,z$ Domain of works (e.g., images, text); individual works
$D$ Dataset $D=(w_{1},\dots,w_{n})$
$\mathsf{Train}$ Training algorithm mapping datasets $D$ to models $p$
$p;~p(y\mid x)$ Conditional generative model; the probability of generating $y$ on prompt $x$
$\mathsf{aux}$ Auxiliary information given to the user (usually $\mathsf{aux}=\mathsf{ideas}(w)$)
$\mathsf{u}$ User taking input $\mathsf{aux}$ and access to $p$ and outputting a work $z\leftarrow\mathsf{u}^{p}(\mathsf{aux})$
$\tau;~\tau(E;\,\mathsf{aux})$ Real-world distribution; the probability of event $z\in E$ in the
probability experiment $p\leftarrow\mathsf{Train}(D)$, $z\leftarrow\mathsf{u}^{p}(\mathsf{aux})$
$\mathcal{C};~c$ Set of in-copyright works; an in-copyright work
$\mathsf{SubSim}(w)$ Set of works substantially similar to $w$
$\mathsf{ideas}(w)$ Non-copyrightable ideas associated with $w$
$w^{\prime}$ stems from $w$ $w^{\prime}$ is a derivative work of $w$, represented as edges in graph $G$
$\mathcal{C}_{D};~\mathcal{C}_{-D}$ In-copyright works from which some work ($\mathcal{C}_{D}$) / no work ($\mathcal{C}_{-D}$) in $D$ stems
$\mathsf{scrub}(D,c)$ Sub-dataset of works in $D$ that don’t stem from $c$
$\tau_{-c};~\tau_{-c}(E;\,\mathsf{aux})$ Clean-room distribution, defined like $\tau$ but replacing $D$ with $\mathsf{scrub}(D,c)$
Table 1: Common symbols and their definitions. Bolded terms are exogenously (legally) defined.
@@ 2 The legal-technical interface
To use legal concepts within a mathematical formalism, we assume some exogenously-defined functions reflecting and operationalizing those concepts. They are the interface between math and law, enabling mathematical study of legal concepts without formalizing the concepts themselves.  Precise definitions of these concepts are out of scope (perhaps impossible, though see [scheffler2022formalizing] for an attempt at formalizing substantial similarity). Three legal concepts are needed for this paper: substantial similarity, ideas, and access. This section defines functions $\mathsf{SubSim}$ and $\mathsf{ideas}$ for the first two. Section 7.1 operationalizes access.
The owner of a copyright in a work has exclusive rights to reproduce, distribute, display, adapt, and perform the work. Original works of authorship are eligible for copyright protections as soon as they are fixed in a tangible medium, and eventually expires. While the work is protected, it is in-copyright. Many types of works are eligible, including books, music, poetry, plays, choreography, photographs, video, and paintings. Copying without permission is sometimes permitted, including under the fair use exception [henderson2023foundation].
@@
Copyright does not protect ideas, only original expression. For example, the text and images in a cookbook may be copyrighted, but not the method of cooking a dish. To be afforded protection, there must be a minimum of creativity beyond the ideas. The function $\mathsf{ideas}$ defines the non-copyrightable ideas of a work, as opposed their expression. We represent these ideas as a string which one can view as encoding a set or any other data type.
@@
We never actually compute $\mathsf{SubSim}$ nor $\mathsf{ideas}$. This is a feature: it means we can be agnostic as to the specific instantiation. One might view them as capturing the “true” legal concept. Or one may take a more practical view. For example, interpreting $\mathsf{SubSim}(w)$ as reflecting the opinion of a typical judge or jury. However the reader wishes to interpret these functions is fine (up to any assumptions stated where used). In particular, our results hold even for $\mathsf{SubSim}(w)=\{w\}$ and $\mathsf{ideas}(w)=\bot$.
In contrast, a model provider (or data provider) would need to actually work with the copyright dependency graph in Section 7. We discuss this limitation in Section 9.
@@ 3 Near access-free models do not provide provable copyright protection
Near access-freeness (NAF) is a mathematical definition for “provable copyright protection” [VyasKB23]. This section describes near access-freeness and its limitations. We prove that models can enable verbatim copying while still satisfying NAF. While NAF provides protection against a single prompt that is independent of the training data, it makes no guarantees against many prompts (composition) [li2024va3], nor a single prompt derived from non-copyrightable ideas (not expression). We emphasize that our focus is definitions for provable copyright protection, where NAF falls short. Practically, NAF may still provide useful protection in many natural settings. For space, we defer algorithms, proofs, and additional discussion to Appendix B.
@@ 3.1 NAF Background
Near access-freeness (NAF) is the first (and only other) attempt at a mathematical definition intended to offer “provable copyright protection” for generative models [VyasKB23]. Motivated by the legal requirement of access, NAF is meant to show “that the defendant’s work is close to a work which was produced without access to the plaintiff’s work” [VyasKB23]. NAF is closely related to DP. But surprisingly, any training algorithm $\mathsf{Train}$ can be used in a black-box way to construct a NAF model without the extra noise that is the hallmark of DP (Theorem 3.3).
@@
Prior works discuss and critique the NAF framework in three ways: questioning NAF’s technical effectiveness [li2024va3], its legal applicability [lee2024talkin, elkinkoren2024copyright], and its real-world practicality [henderson2023foundation, elkinkoren2024copyright, lee2024talkin, chen2024randomization].
$\ast$ Technically, Li, Shen, and Kawaguchi propose and empirically evaluate an attack called VA3 against NAF [li2024va3]. See Appendix B.3 for a full discussion. They show that a black-box attacker, can reliably induce a model to produce outputs that infringe on a target work $c^{*}$. They also give a white-box prompt writing algorithm for diffusion models that greatly improves performance. VA3 provides good evidence that the $\mathsf{CP\text{-}k}$ algorithm of [VyasKB23] may not prevent infringement. Some uncertainty remains because the algorithm implemented in VA3 deviates from the original. The paper leaves open whether $k$-NAF (with fixed $k$) prevents copyright infringement, formalizes a flawed attack model, and avoids the underlying definitional questions almost entirely.
$\ast$ Legally, Elkin-Koren et al. [elkinkoren2024copyright] give the most extensive response to NAF. They correctly argue that copyright cannot be “reduced to privacy.” However, this is a straw man of version of [VyasKB23] and the present paper (see Appendix A). The sharpest criticism is by Lee, Cooper, and Grimmelman who argue that NAF is simply wrong on the law [lee2024talkin]. “It is not a defense to copyright infringement that you would have copied the work from somewhere else if you hadn’t copied it from the plaintiff.” See Appendix B.4 for more discussion of NAF’s legal relevance.
$\ast$ Practically, the main concern is that deduplicating data in the way needed to make NAF effective is infeasible [elkinkoren2024copyright, lee2024talkin, henderson2023foundation]. It is unrealistic to produce enough clean training data to pretrain foundation models, advances in deduplication notwithstanding. Moreover, NAF learning algorithms are computationally expensive and may lag in performance [chen2024randomization]. These concerns are justified and also apply to DP. Still, there may be settings where golden data is practical (Sec. 9), to say nothing of the value of theory.
@@ 3.2 NAF definition and the Copy Protection ($\mathsf{CP}$) algorithm
Near access-freeness (NAF) is defined with respect to a function $\mathsf{safe}$. The $\mathsf{safe}$ function maps a copyrighted data point $c\in\mathcal{C}$ to a generative model $\mathsf{safe}_{c}\in\mathcal{P}$ trained without access to $c$. An example $\mathsf{sharded\mbox{-}safe}$ is in Appendix B (Alg. 1). Thm. 3.3 presents the main feasibility result: any training algorithm $\mathsf{Train}$ can be used as a black-box to construct an NAF model $\mathsf{CP}$, short for copy protection.
@@
Theorem 3.3 (Copy Protection algorithm [VyasKB23]).
@@ 3.3.1 $\mathsf{CP}$ can regurgitate training data
We show that $\mathsf{CP}$—the main NAF algorithm of [VyasKB23]—can fail to protect against copying. Using a prompt containing no copyrightable expression, a user can cause the NAF model returned by $\mathsf{CP}$ to regurgitate copyrighted training data. [^4] This doesn’t violate Theorem 3.3. We use the fact that the $\mathsf{CP}$ algorithm is not actually $k$-NAF for any fixed $k$. It is $k_{x}$-NAF, where $k_{x}$ depends on the prompt $x$. Applied to our construction, the bound on $k_{x}$ given by Theorem 3.3 is not meaningful for $x\in\mathsf{ideas}(D)$. Adapting our construction to the alternate NAF algorithm $\mathsf{CP\text{-}k}$ in [VyasKB23] results in a model sometimes fails to terminate. This is based on an observation of Thomas Steinke [steinkePersonal].
Observe that $\mathsf{CP}$ is a black-box transformation from an underlying training algorithm $\mathsf{Train}$. Thus, $\mathsf{CP}$ is really a family of algorithms—one per $\mathsf{Train}$. The following theorem states that there exists $\mathsf{Train}^{*}$ such that the resulting NAF algorithm $\mathsf{CP}^{*}$ fails in the following way. On prompt $x=\mathsf{ideas}(c)$ reproduces training datum $c\in D$ verbatim, as long as $c$ is uniquely distinguishable by its unprotected ideas.
@@
We defer the proof of (a slight generalization) of this theorem to Appendix B.2. The proof uses the training algorithm $\mathsf{Train}^{*}=\mathsf{Train}_{\mathsf{kv}}$ (Algorithm 3) described in Example 5.2.